OpenStudy (peachytea):

the volume of this circular cylinder is 588pi cm^3. what is the height of this cylinder? the base of it says the radius is 7

OpenStudy (jdoe0001):

\( { \textit{volume of a cylinder}=\pi r^2 h\qquad \begin{cases} r=radius\to 7\\ h=height \end{cases}\qquad 588=\pi \cdot 7^2\cdot h }\) solve for "h' to find the height